OpenStudy (anonymous):

f(x) = 5 – 3x² x = -1 x = 2 Evaluate f(x) at the given values of x.

OpenStudy (lgbasallote):

when x = -1 f(-1) = 5 - 3(-1)^2 f(-1) = 5 - 3(1) f(-1) = 5 - 3 when x = 2 f(2) = 5 - 3(2)^2 f(2) = 5 - 3(4) f(2) = 5 - 12 got it?

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

what about Evaluate the function f(x) at the given values. f(x) = 3x² -6 x = -3 x = 2

OpenStudy (lgbasallote):

just substitute